# Tomographic Alcock-Paczynski Test with Marked Correlation Functions

### v4

@@ III.3 PCA data compression
In principle, utilizing the full shape of the correlation function can maximize the information regarding constraints on cosmology. For instance, the findings of [Park:2019mvn] and [Dong_2023] demonstrate that the $(s,\mu)$-dependent 2-dimensional 2PCF significantly improves compared to the 1-dimensional $\mu$-dependent 2PCF. Consequently, we will also employ 2-dimensional MCFs to constrain cosmology. However, a computational challenge arises in the estimation of the covariance. The accurate estimation of covariance for a large number of statistical quantities depends on a substantial number of mock samples, requiring extensive simulation and computational time.
@@
To address this issue, we propose a principal component analysis (PCA)-based compression scheme that operates directly on the unbinned data, supported by scikit-learn[scikit-learn]. This approach enables efficient $\chi^{2}$ computation while preserving as much information as possible from the full shape of MCFs. PCA reduces the dimensionality of a dataset by transforming correlated variables into an orthogonal set of uncorrelated modes, retaining those with the largest eigenvalues. This minimizes correlations while preserving key information, improving the feasibility and robustness of the analysis. We assume that only the first $N_{c}$ principal components contain relevant cosmological information, with $N_{c}$ determined through simulations based on explained variance ratio $r_{c}$. See details in Sec.IV.2.
It’s a compression algorithm similar to PCA called "massive optimized parameter estimation and data compression" (MOPED, [MOPED]). The algorithm is lossless in theory because it keeps the Fisher information matrices identical. We try to introduce it into our analysis. Although there are some problems and limitations, it still shows great potential. We intend to make a detailed analysis in the future work, while only do preliminary tests and comparisons with PCA in this work. See details in Appendix C.
@@ IV.1 Comparasion between 2PCF and MCFs for the coarse binning scheme
As shown, the overall constraining power of the 2PCF and MCFs is comparable, and none of the individual $\alpha$ values alone significantly reduce the parameter space. However, from the joint $\Omega_{m}$$w$ posterior distributions, we observe that the case with $\alpha=1$ yields notably different constraints compared to the others. Therefore, as pointed out by [Yang:2020ysv], we will combine the 2PCF with MCFs at the four $\alpha$ values in the following analysis and demonstrate that this leads to improved constraints.
@@ IV.3 Impact of redshift errors
In future stage-IV slitless spectroscopic surveys, redshift errors will be non-negligible, reducing resolution along the line of sight and potentially compromising the tomographic AP method [Xiao_2022]. These errors can mimic a false FOG effect, blurring both the AP and RSD signals. Redshift uncertainties distort cosmological measurements by biasing the inferred comoving distances, quantified as $\sigma_{r}=c\sigma_{z}/H(z)$. As a result, clustering information on scales $\lesssim\sigma_{r}$ becomes inaccessible. For instance, the upcoming CSST optical survey expects a redshift error of $\sigma_{z}\sim 0.002$, an order of magnitude larger than current spectroscopic surveys, corresponding to a few Mpc uncertainty in comoving distance.
@@ IV.4 Robustness check about the systematics estimation
A key point of tomographic AP test is the accuracy of systematic estimation. We will not investigate this complex issue in depth here, since the systematic uncertainties depend critically on each survey’s data analysis methodology. Besides, [LI18] found that the change in the cosmological constraints is not significant($\lesssim 0.2\sigma$ for a SDSS-like survey) even ignoring the systematic effect and do not correct it in the joint-analysis with 2PCF.
@@ V Summary and discussion
Our study demonstrates how to effectively extract large-scale structure (LSS) information using the tomographic Alcock-Paczyński (AP) method with MCFs. These techniques can be adapted to other surveys and datasets, potentially helping to constrain additional cosmological parameters. The framework also shows strong potential for future emulator-based analyses with the China Space Station Telescope (CSST). We provide a simple extension of this work to CSST in Appendix D based on the relation of covariance and galaxy number density[LI19]. Besides, according to the number density and survey volume from  [Miao_2022], the constraints could be around 30% tighter than those in this work for a CSST-like survey, without considering other errors. We plan to do a formal preparation in the future, when the mock with accurate systematic estimation of CSST is ready.
In the future, we plan to expand our analysis by considering the methodology developed in  [XiaoYuan_2022], which suggests that the inclusion of the density gradient as a weight may further improve the sensitivity to cosmological parameters. In addition, we do not consider the validation of redshift evolution of RSD in different cosmologies. Some tests have been done, showing that cosmological parameters do have an impact on the RSD evolution but the impact is small [Park:2019mvn]. With the improvement of the accuracy of the stage-IV surveys, it is indeed necessary to consider the validation. We are going to do the validation in the future, when the mock with different $w\mathrm{CDM}$ cosmologies that meet our requirement is ready. We also emphasize that, although this work focuses on MCFs, the tomographic AP method is applicable to almost any kind of statistics, as long as the statistics is capable of probing anisotropic clustering. This makes it an extremely promising tool for exploring nonlinear scales. We plan to continue investigating this approach in future studies.
@@ Appendix A Some tests about insignificant cosmological dependence of redshift evolution of RSD
We first consider it using a physical model for the $\xi(s,\mu;z)$, based on perturbation theory (CLPT[CLPT], EFT[EFT1, EFT2, EFT3], etc.). We adopt the CLPT (Convolution Lagrangian Perturbation Theory), which is a new formulation of Lagrangian perturbation theory that allows accurate predictions of the real space and redshift space correlation functions of the dark matter haloes, and we adopt the Gaussian streaming model [GS_model] to create the redshift space statistics. The code we adopt is CLPT_\_GSRSD [^6] https://github.com/wll745881210/CLPT_\_GSRSD[CLPT_GSRSD]. We fit the first and second order Lagrangian bias parameters: $\langle F^{\prime}\rangle$ and $\langle F^{\prime\prime}\rangle$ with the simulation we use in this work (BIGMD) and then calculate $\xi(s,\mu;z_{1})$ and $\xi(s,\mu;z_{2})$ with RSD.
@@
To confirm this, we then conduct additional simulations based on a fast algorithm: COLA(COmoving Lagrangian Acceleration[COLA, mg_COLA]), which is an N-body method for solving for Large Scale Structure (LSS) in a frame that is comoving with observers following trajectories calculated in Lagrangian Perturbation Theory (LPT). It can straightforwardly trade accuracy at small-scales in order to gain computational speed without sacrificing accuracy at large scales, and according to the tests of [Ma_2020], it meets our requirements in this part.
@@ Appendix C A preliminary test of MOPED
In this section we test the performance of using MOPED rather than PCA to compress the data vectors. MOPED is a compression method which can form linear combinations of the data which contain as much information, in the sense that the Fisher information matrices are identical, if the noise in the data is independent of the parameters; i.e. the method is lossless[MOPED].
@@ Appendix D A simple extension to a CSST-like survey
It would be useful to have a rough estimation of the power of cosmological constraints for a CSST-like survey. In this work, the analysis is done utilizing the BIGMD mock with a length of $2500\mathrm{Mpc}/h$ and fix the density of halos and subhalos with $\overline{n}=10^{-3}$ in the two snapshots of $z=$ 1.0 and 0.6069. We extend our results to a CSST-like survey based on the conclusion in  [LI19]. The authors found that the covariance matrices scale with $N_{\rm gal}$ as
@@
From  [Miao_2022], the number densities in the CSST spectroscopic survey are $5.63\times 10^{-3}$ and $1.15\times 10^{-3}$ in the spec-z bins of $0.6\leq z\leq 0.9$ and $0.9\leq z\leq 1.2$, respectively, which are in the same order of magnitude as the number density of mocks used in our work. The sky coverage of CSST is around 17,500 $\rm deg^{2}$, and therefore, the ratio of comoving volumes between two CSST spec-z bins and BIGMD are around 0.84 and 0.95, respectively. Based on these facts, Eq. 35 tells us that the ratio of variance between CSST and BIGMD is around 0.6. So for CSST, the results of constraints are roughly 30% tighter than those shown in this work.

### v5

@@ III.3 PCA data compression
In principle, utilizing the full shape of the correlation function can maximize the information regarding constraints on cosmology. For instance, the findings of [45] and [17] demonstrate that the $(s,\mu)$-dependent 2-dimensional 2PCF significantly improves compared to the 1-dimensional $\mu$-dependent 2PCF. Consequently, we will also employ 2-dimensional MCFs to constrain cosmology. However, a computational challenge arises in the estimation of the covariance. The accurate estimation of covariance for a large number of statistical quantities depends on a substantial number of mock samples, requiring extensive simulation and computational time.
@@
To address this issue, we propose a principal component analysis (PCA)-based compression scheme that operates directly on the unbinned data, supported by scikit-learn[46]. This approach enables efficient $\chi^{2}$ computation while preserving as much information as possible from the full shape of MCFs. PCA reduces the dimensionality of a dataset by transforming correlated variables into an orthogonal set of uncorrelated modes, retaining those with the largest eigenvalues. This minimizes correlations while preserving key information, improving the feasibility and robustness of the analysis. We assume that only the first $N_{c}$ principal components contain relevant cosmological information, with $N_{c}$ determined through simulations based on explained variance ratio $r_{c}$. See details in Sec.IV.2.
It’s a compression algorithm similar to PCA called "massive optimized parameter estimation and data compression" (MOPED, [24]). The algorithm is lossless in theory because it keeps the Fisher information matrices identical. We try to introduce it into our analysis. Although there are some problems and limitations, it still shows great potential. We intend to make a detailed analysis in the future work, while only do preliminary tests and comparisons with PCA in this work. See details in Appendix C.
@@ IV.1 Comparasion between 2PCF and MCFs for the coarse binning scheme
As shown, the overall constraining power of the 2PCF and MCFs is comparable, and none of the individual $\alpha$ values alone significantly reduce the parameter space. However, from the joint $\Omega_{m}$$w$ posterior distributions, we observe that the case with $\alpha=1$ yields notably different constraints compared to the others. Therefore, as pointed out by [66], we will combine the 2PCF with MCFs at the four $\alpha$ values in the following analysis and demonstrate that this leads to improved constraints.
@@ IV.3 Impact of redshift errors
In future stage-IV slitless spectroscopic surveys, redshift errors will be non-negligible, reducing resolution along the line of sight and potentially compromising the tomographic AP method [64]. These errors can mimic a false FOG effect, blurring both the AP and RSD signals. Redshift uncertainties distort cosmological measurements by biasing the inferred comoving distances, quantified as $\sigma_{r}=c\sigma_{z}/H(z)$. As a result, clustering information on scales $\lesssim\sigma_{r}$ becomes inaccessible. For instance, the upcoming CSST optical survey expects a redshift error of $\sigma_{z}\sim 0.002$, an order of magnitude larger than current spectroscopic surveys, corresponding to a few Mpc uncertainty in comoving distance.
@@ IV.4 Robustness check about the systematics estimation
A key point of tomographic AP test is the accuracy of systematic estimation. We will not investigate this complex issue in depth here, since the systematic uncertainties depend critically on each survey’s data analysis methodology. Besides, [36] found that the change in the cosmological constraints is not significant($\lesssim 0.2\sigma$ for a SDSS-like survey) even ignoring the systematic effect and do not correct it in the joint-analysis with 2PCF.
@@ V Summary and discussion
Our study demonstrates how to effectively extract large-scale structure (LSS) information using the tomographic Alcock-Paczyński (AP) method with MCFs. These techniques can be adapted to other surveys and datasets, potentially helping to constrain additional cosmological parameters. The framework also shows strong potential for future emulator-based analyses with the China Space Station Telescope (CSST). We provide a simple extension of this work to CSST in Appendix D based on the relation of covariance and galaxy number density[33]. Besides, according to the number density and survey volume from  [44], the constraints could be around 30% tighter than those in this work for a CSST-like survey, without considering other errors. We plan to do a formal preparation in the future, when the mock with accurate systematic estimation of CSST is ready.
In the future, we plan to expand our analysis by considering the methodology developed in  [65], which suggests that the inclusion of the density gradient as a weight may further improve the sensitivity to cosmological parameters. In addition, we do not consider the validation of redshift evolution of RSD in different cosmologies. Some tests have been done, showing that cosmological parameters do have an impact on the RSD evolution but the impact is small [45]. With the improvement of the accuracy of the stage-IV surveys, it is indeed necessary to consider the validation. We are going to do the validation in the future, when the mock with different $w\mathrm{CDM}$ cosmologies that meet our requirement is ready. We also emphasize that, although this work focuses on MCFs, the tomographic AP method is applicable to almost any kind of statistics, as long as the statistics is capable of probing anisotropic clustering. This makes it an extremely promising tool for exploring nonlinear scales. We plan to continue investigating this approach in future studies.
@@ Appendix A Some tests about insignificant cosmological dependence of redshift evolution of RSD
We first consider it using a physical model for the $\xi(s,\mu;z)$, based on perturbation theory (CLPT[12], EFT[25, 11, 51], etc.). We adopt the CLPT (Convolution Lagrangian Perturbation Theory), which is a new formulation of Lagrangian perturbation theory that allows accurate predictions of the real space and redshift space correlation functions of the dark matter haloes, and we adopt the Gaussian streaming model [52] to create the redshift space statistics. The code we adopt is CLPT_\_GSRSD [^6] https://github.com/wll745881210/CLPT_\_GSRSD[59]. We fit the first and second order Lagrangian bias parameters: $\langle F^{\prime}\rangle$ and $\langle F^{\prime\prime}\rangle$ with the simulation we use in this work (BIGMD) and then calculate $\xi(s,\mu;z_{1})$ and $\xi(s,\mu;z_{2})$ with RSD.
@@
To confirm this, we then conduct additional simulations based on a fast algorithm: COLA(COmoving Lagrangian Acceleration[58, 63]), which is an N-body method for solving for Large Scale Structure (LSS) in a frame that is comoving with observers following trajectories calculated in Lagrangian Perturbation Theory (LPT). It can straightforwardly trade accuracy at small-scales in order to gain computational speed without sacrificing accuracy at large scales, and according to the tests of [40], it meets our requirements in this part.
@@ Appendix C A preliminary test of MOPED
In this section we test the performance of using MOPED rather than PCA to compress the data vectors. MOPED is a compression method which can form linear combinations of the data which contain as much information, in the sense that the Fisher information matrices are identical, if the noise in the data is independent of the parameters; i.e. the method is lossless[24].
@@ Appendix D A simple extension to a CSST-like survey
It would be useful to have a rough estimation of the power of cosmological constraints for a CSST-like survey. In this work, the analysis is done utilizing the BIGMD mock with a length of $2500\mathrm{Mpc}/h$ and fix the density of halos and subhalos with $\overline{n}=10^{-3}$ in the two snapshots of $z=$ 1.0 and 0.6069. We extend our results to a CSST-like survey based on the conclusion in  [33]. The authors found that the covariance matrices scale with $N_{\rm gal}$ as
@@
From  [44], the number densities in the CSST spectroscopic survey are $5.63\times 10^{-3}$ and $1.15\times 10^{-3}$ in the spec-z bins of $0.6\leq z\leq 0.9$ and $0.9\leq z\leq 1.2$, respectively, which are in the same order of magnitude as the number density of mocks used in our work. The sky coverage of CSST is around 17,500 $\rm deg^{2}$, and therefore, the ratio of comoving volumes between two CSST spec-z bins and BIGMD are around 0.84 and 0.95, respectively. Based on these facts, Eq. 35 tells us that the ratio of variance between CSST and BIGMD is around 0.6. So for CSST, the results of constraints are roughly 30% tighter than those shown in this work.